# Approximating Two-Layer ReLU Networks for Hidden State Analysis in Differential Privacy

### v2

@@ 6 Experimental Results
We compare the methods on standard benchmark image datasets: MNIST (LeCun et al.,, 1998), FashionMNIST (Xiao et al.,, 2017) and CIFAR-10 (Krizhevsky et al.,, 2009). The MNIST datasets have a training dataset of $6\cdot 10^{4}$ samples and a test dataset of $10^{4}$ samples and CIFAR-10 has a training dataset size of $5\cdot 10^{4}$ and a test dataset of $10^{4}$ samples. All samples in MNIST datasets are $28\times 28$ size gray-level images and in CIFAR-10 $32\times 32$ color images (with three channels each). We experimentally compare three alternatives: DP-SGD applied to a one-hidden layer ReLU network, DP-SGD applied to the stochastic convex model (3.6) without regularization ($\lambda=0$) and NoisyCGD applied to the stochastic convex model (3.6) with $\lambda>0$. We use the cross-entropy loss for all the models considered. In order to simplify the comparisons, we fix the batch size to 1000 for all the methods and train all methods for 400 epochs. We compare the results on two noise levels: $\sigma=5.0$ and $\sigma=15.0$.
Although one hidden-layer networks with tempered sigmoid activations (Papernot et al.,, 2021) would likely yield improved results, we focus on ReLU networks as baselines for consistency as this allows us to compare the methods in the context of ReLU-based architectures. Also, this does not affect our main finding that we are able to find improved models for the hidden state analysis.
@@
Unlike in the experiments of, e.g., (Abadi et al.,, 2016), we do not use pre-trained convolutive layers for obtaining higher test accuracies in the CIFAR-10 experiment, as we experimentally observe that the DP-SGD trained logistic regression gives similar accuracies as the DP-SGD trained ReLU network. Thus, we consider the much more difficult problem of training the models from scratch using the vectorized CIFAR-10 images as features.
The hyperparameter tuning of NoisyCGD is simplified by the fact that bound of Thm. (2.8) depends monotonoysly on the parameter $c=1-\eta\cdot\lambda$. In case the hyperparameters $b$ noise scale $\sigma$ are fixed, fixing the GDP parameter $\mu$ will also fix the value of $c$. Thus, if we have a grid of learning rate candidates those will also determine the values of $\lambda$’s as well. Overall, in case the batch size, number of epochs and $\sigma$ are fixed, in addition to the learning rate $\eta$, we have in all alternatives only one hyperparameter to tune: the hidden-layer width $W$ for the ReLU networks and the number of hyperplanes $P$ for the convex models. The hyperparameter grids used in the experiments are depicted in Appendix H.
We generally find that $P=128$ is not far from the optimum for the convex model (see Appendix E.2 for comparisons using MNIST). Appendix E.1 also shows that in the non-private case, the stochastic problem approximates the ReLU problem better as $P$ increases.
6.1 Results on DP Hyperparameter Tuning
Tables 1 and 2 show the accuracies of the best models obtained using the DP hyperparameter tuning algorithm depicted in Section 5. With the noise scale values $\sigma=5.0$ and $\sigma=15.0$, the DP-SGD trained models are $(1.33,10^{-5})$-DP and $(4.76,10^{-5})$-DP, respectively. To have similar privacy guarantees for the base models trained using NoisyCGD, we adjust the regularization constant $\lambda$ accordingly (as depicted in Secion 5) which leads to equal final $(\varepsilon,\delta)$-DP guarantees for the hyperparameter tuning algorithms. We see from tables 1 and 2 that the convex models are on par in accuracy with the ReLU network. Notice also from the results of Section 6.2 that the learning rate tuned logistic regression cannot reach similar accuracies as the NoisyCGD trained convex model.
Table 1: MNIST model accuracies vs. $\varepsilon$-values for the DP hyperparameter tuning algorithm when $\delta=10^{-5}$. The number of candidate models $K$ is Poisson distributed with mean $20$. Results are means of 5 Runs.
$\varepsilon$ NoisyCGD + Convex Approx. DP-SGD + Convex Approx. DP-SGD + ReLU
2.88 0.927 0.929 0.916
8.91 0.944 0.949 0.944
Table 2: CIFAR-10 model accuracies vs. $\varepsilon$-values for the DP hyperparameter tuning algorithm when $\delta=10^{-5}$. The number of candidate models $K$ is Poisson distributed with mean $m=20$. Results are means of 5 Runs.
$\varepsilon$ NoisyCGD + Convex Approx. DP-SGD + Convex Approx. DP-SGD + ReLU
2.88 0.416 0.417 0.427
8.91 0.455 0.459 0.471
6.2 Result with the Best Models
@@
Figures 2 and 3 show the accuracies of the hyperparameter tuned models along the the training iteration of 400 epochs for the MNIST and CIFAR-10 experiment, when the privacy cost of the tuning is not taken into account. We give results for the FashionMNIST experiment in Appendix I. Figures 2 and 3 include additionally the accuracies for the learning rate optimized logistic regression models. We observe that the proposed convex model significantly outperforms logistic regression, which has been the most accurate model considered in the literature for hidden state DP analysis up to now.
@@
Figure 2 show that the convexification helps in the MNIST experiment: both DP-SGD and the noisy cyclic mini-batch GD applied to the stochastic dual problem lead to better utility models than DP-SGD applied to the ReLU network. Notice also that the final accuracies for DP-SGD are not far from the accuracies obtained by Abadi et al., (2016) using a three-layer network for corresponding $\varepsilon$-values which can be compared using the fact that there is approximately a multiplicative difference of 2 between the two relations: the add/remove neighborhood relation used by (Abadi et al.,, 2016) and the substitute neighborhood relation used in our work.
@@
Results of figures 23 and I and are averaged over 5 trials and the error bars on both sides of the mean values depict 1.96 times the standard error.
Figure 2: MNIST Comparisons: Test accuracies vs. the spent privacy budget $\varepsilon$, when $\delta=10^{-5}$ and each model is trained for 400 epochs. The ReLU network is a one hidden-layer fully connected network, and the batch size equals 1000 for all methods considered.
Figure 3: Cifar-10 Comparisons: Test accuracies vs. the spent privacy budget $\varepsilon$, when $\delta=10^{-5}$ and each model is trained for 400 epochs. The model is a one hidden-layer fully connected ReLU network and the batch size equals 1000 for all methods considered.
@@ 7 Conclusions and Outlook
We have shown how to privately approximate the two-layer ReLU network and we have given the first high privacy-utility trade-off results using the hidden state DP analysis. In particular, we have given the first high privacy-utility trade-off results for the noisy cyclic mini-batch GD which makes it more suitable for practical applications of DP ML model training. As shown by our experiments on benchmark image classification datasets, the results for the convex problems have similar privacy-utility trade-offs as those obtained by applying DP-SGD to a one hidden-layer ReLU network and using the composition analysis. Theoretically, an interesting future task is to carry out end-to-end utility analysis for private optimization of ReLU networks via the dual form. The recent results by Kim and Pilanci, (2024) could be helpful for this as they show connections between the stochastic approximation of the dual form and the ReLU minimization problem. Also, an interesting general question is, whether it is possible to obtain still better privacy-utility trade-offs for the final model in the hidden state threat model by using the privacy amplification by iteration type of analysis of, for example, DP-SGD. In order to get a better understanding of this question, tighter privacy amplification by iteration analysis for, e.g., DP-SGD or Noisy-CGD would be needed, as the composition analysis of DP-SGD cannot likely be improved a lot. Furthermore, developing DP convex models that approximate deeper neural networks, including those with convolutional layers, is an intriguing direction for future research.
@@ Appendix A Appendix
Appendix B Reference Algorithm for the Utility Bounds
Algorithm 1 Differentially Private Gradient Descent (Song et al.,, 2021)
1:  Input: dataset $D=\{D_{i}\}_{i=1}^{n}$, loss function $\ell\,:\,\mathbb{R}^{p}\times\mathcal{X}\rightarrow\mathbb{R}$, gradient $\ell_{2}$-norm bound $L$, constraint set $\mathcal{C}\subseteq\mathbb{R}^{p}$, number of iterations $T$, noise variance $\sigma^{2}$, learning rate $\eta$.
@@ Appendix B Reference Algorithm for the Utility Bounds
4:     $g^{\textrm{priv}}_{t}\leftarrow\frac{1}{n}\sum\limits_{i=1}^{n}\partial_{\theta}\ell(\theta_{t},d_{i})+b_{t}$, where $b_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$.
@@
Appendix C Proof of Lemma 3.2
Lemma C.1.
The gradients of the loss function
$\ell(v,x_{j},y_{j})=\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}(D_{i})_{jj}x_{j}^{T}v_{i}-y_{j}\right\rVert_{2}^{2}+\frac{\lambda}{2}\sum\nolimits_{i=1}^{P}\left\lVert v_{i}\right\rVert_{2}^{2}$
Proof.
For the quadratic function
$h(v)=\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}(D_{i})_{jj}x_{j}^{T}v_{i}-y_{j}\right\rVert_{2}^{2}$
the Hessian matrix is of the diagonal block form
$\nabla^{2}h=\mathrm{diag}\left(A_{1},\ldots,A_{P}\right),$
where $A_{i}=x_{j}D_{i}^{2}x_{j}^{T}=x_{j}D_{i}x_{j}^{T}$. Since for all $i\in[P]$, $x_{j}D_{i}x_{j}^{T}\preccurlyeq x_{j}x_{j}^{T}$, $\nabla^{2}h\preccurlyeq x_{j}x_{j}^{T}$ and furthermore for the spectral norm of $\nabla^{2}h$ we have that $\left\lVert\nabla^{2}h\right\rVert_{2}\leq\left\lVert x_{j}x_{j}^{T}\right\rVert_{2}=\left\lVert x_{j}\right\rVert_{2}^{2}$. ∎
Appendix D Utility Bound within the Random Data Model
We give a convergence analysis for the minimization of the loss function
$h(v)=\frac{1}{2}\left\lVert\sum\limits_{i=1}^{P}D_{i}Xv_{i}-y\right\rVert_{2}^{2}$
without any constraints. Kim and Pilanci, (2024) have recently given several results for the stochastic approximations of the convex problem (3.3). The analysis uses the condition number $\kappa$ defined as
$\kappa=\frac{\lambda_{\max}(XX^{T})}{\lambda_{\min}(M)},$

### v3

@@ 5 Experimental Results
We compare the methods on standard image classification benchmarks: MNIST (LeCun et al., , 1998), FashionMNIST (Xiao et al., , 2017), and CIFAR10 (Krizhevsky et al., , 2009). The MNIST dataset has a training dataset of $6\cdot 10^{4}$ samples and a test dataset of $10^{4}$ samples. CIFAR10 has a training dataset size of $5\cdot 10^{4}$ and a test dataset of $10^{4}$ samples. All samples in MNIST datasets are $28\times 28$ size gray-level images, and in CIFAR10 $32\times 32$ color images (with three color channels each). We compare three alternatives: DP-SGD applied to a 2-layer ReLU network, DP-SGD applied to the stochastic convex model (3.6) without regularization ($\lambda=0$), and NoisyCGD applied to the stochastic convex model (3.6) with $\lambda>0$. We use the cross-entropy loss for all the models considered. To simplify the comparisons, we set the batch size to 1000 for all methods and train them for 400 epochs. We compare the results on two noise levels, $\sigma=5.0$ and $\sigma=15.0$, which correspond to privacy parameter $\varepsilon=1.33$ and $\varepsilon=4.76$ at $\delta=10^{-5}$. In all experiments, the model parameters are always initialized to zero.
@@
Although 2-layer networks with tempered sigmoid activations (Papernot et al., , 2021) would likely yield improved results, we focus on ReLU networks as baselines for consistency. This allows us to compare the methods in the context of ReLU-based architectures and will not affect our main finding, which is that we can find improved models for the hidden state privacy analysis. Unlike Abadi et al., (2016), we do not use pre-trained convolutive layers for obtaining higher test accuracies in the CIFAR10 experiment, as we experimentally observe that the DP-SGD trained logistic regression gives similar accuracies as the DP-SGD trained ReLU network. We consider the much more difficult problem of training the models from scratch using the vectorized CIFAR10 images as features.
We also compare the hyperparameter tuning under differential privacy and describe this in Appendix F. The experimental results in this section use the hyperparameter from the private tuning process. The hyperparameter tuning of NoisyCGD is simplified by the fact that the bound of Thm. (2.7) depends monotonously on the parameter $c=1-\eta\cdot\lambda$. If the hyperparameters $b$ and noise scale $\sigma$ are fixed, fixing the GDP parameter $\mu$ also fixes the value of $c$. Thus, a grid of learning rate candidates also determines the values of $\lambda$’s. The hyperparameter grids are noted in Appendix F.1.
Baselines. As the stochastic approximation yields a convex problem, we compare, as a baseline, to a linear regression model that uses the same features. In contrast to the iterative methods, we use the Sufficient Statistics Perturbation method, which is a one-step solution (SSP Amin et al., , 2023). SSP is a state of the art model for linear regression with Differential Privacy guarantees, and the details are provided in Appendix G. We consider three sets of features: a) the stochastic approximation features in the GLM formulation of Equation 3.7 (Convex approx.), b) the activation patterns of Equation 3.3 (Random ReLU), and c) Random Fourier Features of the input images Rahimi and Recht, (2007).
Table 1: A comparison of model accuracies vs. $\varepsilon$-values. The iterative methods generally score better than SSP and have comparable accuracy, which shows a high-utility result for hidden-state privacy analysis of a 2-layer neural network. The results are the mean accuracy among five random restarts.
MNIST CIFAR-10
$\varepsilon=1.33$ $\varepsilon=4.76$ $\varepsilon=1.33$ $\varepsilon=4.76$
Sufficient Statistics Perturbation (Convex Approx.) $51.9_{{\color[rgb]{.5,.5,.5}\pm 1.1}}$ $67.0_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $19.2_{{\color[rgb]{.5,.5,.5}\pm 1.1}}$ $23.3_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$
Sufficient Statistics Perturbation (Random ReLU) $52.5_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $65.6_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $19.3_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $25.9_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$
Sufficient Statistics Perturbation (RFF) $64.1_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $77.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $21.3_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $28.5_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$
DP-SGD + Convex Approximation $93.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $94.9_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $41.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $45.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$
DP-SGD + ReLU $91.7_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $94.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $42.5_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $47.0_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$
NoisyCGD + Convex Approximation $92.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $94.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $41.0_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $45.4_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$
@@
Main Results. Figures 2 and 3 show the accuracies of the models along a training of 400 epochs for the MNIST and CIFAR10 dataset. The results for FashionMNIST are in Appendix H. The figures also include the accuracies for the learning rate-optimized logistic regression models. We observe that the proposed convex model significantly outperforms logistic regression, which has been the most accurate model considered in the literature for hidden state DP analysis. The hyperparameters for this experiment were found with Differentially Private hyperparameter tuning (Appendix F).
@@
Figure 2 shows that the convexification helps in the MNIST experiment: both DP-SGD and the noisy cyclic mini-batch GD applied to the stochastic dual problem lead to better utility models than DP-SGD applied to the ReLU network. Notice also that the final accuracies for DP-SGD are not far from the accuracies obtained by Abadi et al., (2016) using a three-layer network for corresponding $\varepsilon$-values which can be compared using the fact that there is approximately a multiplicative difference of 2 between the two relations: the add/remove neighborhood relation used by (Abadi et al., , 2016) and the substitute neighborhood relation used in our work.
@@
The results of Figures 23 and Section H are averaged over five trials. The error bars on both sides of the mean values indicate 1.96 times the standard error. In all experiments, we set $\delta=10^{-5}$.
The proposed method is also compared to non-iterative baseline methods. Table 1 compares NoisyCGD and DP-SGD against Sufficient Statistics Perturbation on random and non-linear features. For each dataset and privacy setting, even the best linear regression model has lower accuracy than any of the iterative methods. This shows that iterative methods for non-linear models can achieve higher accuracy, and we believe that NoisyCGD is an important step in this direction.
Figure 2: MNIST: Test accuracy versus the spent privacy budget $\varepsilon$, when each model is trained for 400 epochs. NoisyCGD and DP-SGD generally have comparable performance for the 2-layer ReLU network and much higher accuracy than logistic regression.
Figure 3: CIFAR10  Test accuracy versus the spent privacy budget $\varepsilon$, when each model is trained for 400 epochs. NoisyCGD and DP-SGD generally have comparable performance for the 2-layer ReLU network and much higher accuracy than logistic regression.
@@ 6 Conclusions and Outlook
We have shown how to privately approximate the two-layer ReLU network, and we have given the first high privacy-utility trade-off results using the hidden state privacy analysis. In particular, we have provided results for the noisy cyclic mini-batch GD, which is more suitable for practical applications of private ML model training than variants of DP-SGD that carry out random subsampling at each iteration. Experimentally, on benchmark image classification datasets, the results for the convex problems have similar privacy-utility trade-offs as those obtained by applying DP-SGD for a 2-layer ReLU network and using the composition analysis. An interesting task for future research is to obtain hidden state privacy guarantees for ‘deeper’ neural networks, for example, by developing convex models that approximate deeper neural networks, including those with convolutional layers (Ergen and Pilanci, , 2020) and different activation functions (Ergen et al., , 2023).
@@ Appendix A Appendix
In this appendix, we provide detailed proofs and additional results that complement the main text. Section B clarifies the connection between the proposed strongly convex models and generalized linear models. In Section C, we prove that the gradients of our loss function are Lipschitz continuous. In Section D, we provide a convergence analysis to minimize the loss function within the random data model, with Section D.1 presenting the reference algorithm for the utility bounds. Appendix E discusses the approximability of the stochastic approximation for the dual problem in the non-private case. Sections F and G contain details on the hyperparameter tuning and experimental setting, respectively, and the SectionH gives complementary results on the FMNIST dataset. Section I illustrates why the shuffling amplification is a poor alternative for obtaining DP guarantees for DP-SGD with disjoint mini-batches, further motivating our contributions. Finally, Section J provides a summary of the notation used throughout the paper.
Appendix B Formulating the Strongly Convex Approximation as a GLM
We first show that the strongly convex loss function given in Eq. (3.6) corresponds to a loss function of a convex generalized linear model. The loss function in Eq. (3.6) is of the form
$\mathcal{L}\big{(}v,X,y\big{)}=\frac{1}{n}\sum\nolimits_{j=1}^{n}\ell_{j}(v,x_{j},y_{j}),$
where
$\ell_{j}(v,x_{j},y_{j})=\frac{1}{2}\left\lVert\sum_{i=1}^{P}(\Lambda_{i})_{jj}x_{j}^{T}v_{i}-y_{j}\right\rVert_{2}^{2}+\frac{\lambda}{2}\sum_{i=1}^{P}\left\lVert v_{i}\right\rVert_{2}^{2},\quad(\Lambda_{i})_{jj}=\mathds{1}(x_{j}^{T}u_{i}\geq 0).$
The $u_{i}$’s are the randomly sampled vectors that determine the boolean matrices $\Lambda_{i}$’s (and the functions $\ell_{j}$), and we denote the vector
$v=\begin{bmatrix}v_{1}\\ \vdots\\ v_{P}\end{bmatrix}\in\mathbb{R}^{P\cdot d}.$
This is actually a generalized linear model: if we denote
$\tilde{x}_{j}=\begin{bmatrix}(\Lambda_{1})_{jj}x_{j}\\ \vdots\\ (\Lambda_{P})_{jj}x_{j}\end{bmatrix},$
we see that
$\ell_{j}(v,x_{j},y_{j})=\frac{1}{2}\left\lVert\tilde{x}_{j}^{T}v-y_{j}\right\rVert_{2}^{2}+\frac{\lambda}{2}\left\lVert v\right\rVert_{2}^{2},$
which shows that we are actually minimizing a loss function of a GLM when we are minimizing the loss $\mathcal{L}\big{(}v,X,y\big{)}$ w.r.t. $v$. By the results of (Song et al., , 2021), we know that the clipped gradients are gradients of an auxiliary convex loss, which allows using the privacy amplification by iteration analysis by Bok et al., (2024).
Moreover, the convexity properties of the GLM loss function are preserved under gradient clipping. This is shown in Appendix E.2 of (Redberg et al., , 2024). In summary, we can use the convexity properties shown in our Section 3.4 for the privacy analysis.
Appendix C Proof of Lemma 3.1
Lemma C.1.
The gradients of the loss function
$\ell(v,x_{j},y_{j})=\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}(\Lambda_{i})_{jj}x_{j}^{T}v_{i}-y_{j}\right\rVert_{2}^{2}+\frac{\lambda}{2}\sum\nolimits_{i=1}^{P}\left\lVert v_{i}\right\rVert_{2}^{2}$
Proof.
For the quadratic function
$h(v)=\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}(\Lambda_{i})_{jj}x_{j}^{T}v_{i}-y_{j}\right\rVert_{2}^{2}$
the Hessian matrix is of the diagonal block form
$\nabla^{2}h=\mathrm{diag}\left(A_{1},\ldots,A_{P}\right),$
where $A_{i}=x_{j}\Lambda_{i}^{2}x_{j}^{T}=x_{j}\Lambda_{i}x_{j}^{T}$. Since for all $i\in[P]$, $x_{j}\Lambda_{i}x_{j}^{T}\preccurlyeq x_{j}x_{j}^{T}$, $\nabla^{2}h\preccurlyeq x_{j}x_{j}^{T}$ and furthermore for the spectral norm of $\nabla^{2}h$ we have that $\left\lVert\nabla^{2}h\right\rVert_{2}\leq\left\lVert x_{j}x_{j}^{T}\right\rVert_{2}=\left\lVert x_{j}\right\rVert_{2}^{2}$. ∎
Appendix D Utility Bound within the Random Data Model
D.1 Reference Algorithm for the Utility Bounds
We consider the following form of DP gradient descent (DP-GD) for the theoretical utility analysis.
Algorithm 1 Differentially Private Gradient Descent (Repeated from Song et al., , 2021)
1:  Input: dataset $D=\{z_{i}\}_{i=1}^{n}$, loss function $\ell\,:\,\mathbb{R}^{p}\times\mathcal{X}\rightarrow\mathbb{R}$, gradient $\ell_{2}$-norm bound $L$, constraint set $\mathcal{C}\subseteq\mathbb{R}^{p}$, number of iterations $T$, noise variance $\sigma^{2}$, learning rate $\eta$.
@@ D.1 Reference Algorithm for the Utility Bounds
4:     $g^{\textrm{priv}}_{t}\leftarrow\frac{1}{n}\sum\limits_{i=1}^{n}\partial_{\theta}\ell(\theta_{t},z_{i})+b_{t}$, where $b_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$.
@@
We have the following classical result for Alg. 1.
Theorem D.1 (Bassily et al., 2014; Talwar et al., 2014).
D.2 Utility Analysis
We give a convergence analysis for the minimization of the loss function (3.5) with $\lambda=0$, i.e., our minimization problem is
$\min_{v}h(v)=\frac{1}{2}\left\lVert\sum\limits_{i=1}^{P}\Lambda_{i}Xv_{i}-y\right\rVert_{2}^{2},$ (D.1)
where the diagonal boolean matrices $\Lambda_{1},\ldots,\Lambda_{P}\in\mathbb{R}^{n\times n}$ are constructed by taking first $P$ i.i.d. samples $u_{1},\ldots,u_{P}$, $u_{i}\sim\mathcal{N}(0,I_{d})$, and then setting the diagonals $\Lambda_{i}$’s: $(\Lambda_{i})_{jj}=\max\big{(}0,\mathrm{sign}(x_{j}^{T}u_{i})\big{)}.$
Kim and Pilanci, (2024) have recently given several results for the minimization problem (3.4). The analysis uses the condition number $\kappa$ defined as
$\kappa=\frac{\lambda_{\max}(XX^{T})}{\lambda_{\min}(\Sigma)},$